### v3

# How Multi-Modal LLMs Reshape Visual Deep Learning Testing? A Comprehensive Study Through the Lens of Image Mutation

Abstract
Visual deep learning (VDL) systems have demonstrated their capability to understand complex image semantics, paving the way for significant real-world applications like image recognition, object detection, and autonomous driving. To evaluate the reliability of VDL, a mainstream approach is software testing, which requires diverse mutations over image semantics. The rapid development of multi-modal large language models (MLLMs) has introduced revolutionary image mutation potentials through instruction-driven methods. Users can now freely describe desired mutations and let MLLMs generate the mutated images. Hence, parallel to large language models’ (LLMs) recent success in traditional software fuzzing, one may also expect MLLMs to be promising for VDL testing in terms of offering unified, diverse, and complex image mutations.
However, the quality and applicability of MLLM-based mutations in VDL testing remain largely unexplored. We present the first study, aiming to assess MLLMs’ adequacy from 1) the semantic validity of MLLMs’ mutated images, 2) the alignment of MLLMs’ mutated images with their text instructions (prompts), and 3) the faithfulness of how different mutations preserve semantics that are ought to remain unchanged. With large-scale human studies and quantitative analysis, we identify MLLM’s promising potential in expanding the covered semantics of image mutations. Notably, while SoTA MLLMs (e.g., GPT-4V) fail to support or perform worse in editing existing semantics in images (as in traditional mutations like rotation), they generate high-quality test inputs using “semantic-replacement” mutations (e.g., “dress a dog with clothes”), which bring extra semantics to images and are effective in triggering VDL faults; these were infeasible for past approaches. Hence, we view MLLM-based mutations as a vital complement to traditional mutations, and advocate future VDL testing tasks to combine MLLM-based and traditional image mutations for comprehensive and reliable testing.
Index Terms:
Deep learning testing, metamorphic testing, multi-modal large language models
1 Introduction
Visual deep learning (VDL) systems comprehend the semantics of image contents to enable a wide range of real-world applications (e.g., autonomous driving, medical imaging diagnosis) that were previously human-reliant. However, VDL systems can experience drastic failures in their outputs due to their obscure decision rules, jeopardizing the reliability of VDL systems in safety-critical applications. Software testing has been very effective in detecting VDL failures; it mutates input images and defines the accompanying testing oracle and numeric validation metric. With the metric validating the achieved mutation, a VDL failure can be detected as a violation of the oracle. In recent years, our community has proposed a variety of mutation schemes for VDL testing under different scenarios, such as image classification [1], object detection [2], image captioning [3], and autonomous driving [4, 5, 6].
Inputs of VDL systems are images whose contents have rich semantics. Thus, to evaluate the reliability of VDL systems via software testing, comprehensively covering different image semantics is demanding for the input mutation. Nevertheless, mutating image semantics is fundamentally challenging because the relation between pixel values and most image semantics is unclear and hard to model [7, 8, 9, 10, 11, 12]; most existing methods primarily mutate some pixel-level semantics such as brightness and contrast [4]. Some recent works support mutating more complex semantics via representation learning techniques [13, 6, 9], which however require paired images containing the expected mutations and only apply to specific domains such as driving scenes [6] or face photos [9].
The recent multi-modal large language models (MLLMs) [^1] Since MLLMs are often jointly used with other large models such as Diffusion model [14], CLIP [15], DALLE [16], etc., to minimize confusion, we use MLLMs as a general term for them. We use their own names when referring to their special usage., e.g., GPT-4V [17], have liberated the freedom of mutations for image semantics: users can freely describe the anticipated mutation in natural language and let MLLMs generate the mutated images. This new instruction-driven mutation paradigm is shedding light on (1) supporting different traditional mutations in a unified prompt-driven manner, and therefore, one may expect to easily understand, maintain, extend, and disseminate various input mutation methods in natural language texts. (2) Moreover, it also shows promising potential in offering a wide range of new mutations that are hardly achievable by traditional methods, based on the high natural language comprehension ability of LLMs, and high image synthesis capabilities of advanced vision backends (e.g., Diffusion model [14]). Thus, one may reasonably anticipate that MLLMs can revolutionize VDL testing, in parallel to large language model (LLM)’s emerging adoption and success in our community to boost traditional software testing [18, 19, 20], program repairing [21, 22, 23], and program synthesis [24, 25, 26].
Nevertheless, unlike traditional methods that explicitly form the mutations (e.g., via mathematical formulas; see Sec. 3.1), it is inherently challenging to interpret how MLLMs generate mutated outputs. Some recent reports show that MLLMs can generate semantically invalid images [27, 28, 29], which makes the follow-up testing meaningless, as those detected “failures” are induced by the VDL system’s incorrect usages, not its faults. Thus, it is demanding to systematically evaluate MLLM’s capability in mutating images.
Following common practice in VDL testing [5, 6, 1], we first measure the quality of MLLM-generated test inputs from: 1$\rangle$ the validity of the overall semantics in mutated images, 2$\rangle$ the alignment of mutated images with text instructions, and 3$\rangle$ the faithfulness of how different mutations preserve semantics that should be unchanged. This is achieved via a large-scale human evaluation on Amazon Mechanical Turk [30], with 20 Ph.D. students experienced in VDL systems and software testing. Based on the quality assessment, we then study MLLM’s mutation capabilities by answering two research questions (RQs). RQ1: Can MLLMs unify different traditional input mutations? RQ2: What benefits can MLLMs bring to input mutations in VDL testing? Since input validation is crucial to VDL testing, we further study if existing validation metrics apply to MLLM-based mutations (RQ3).
We consider two recently proposed (the only two to our knowledge) pipelines of leveraging MLLMs for image mutations; both of them implement optimizations for prompts and image generation modules to provide out-of-the-box usages. Our evaluations incorporate four SoTA MLLMs (including GPT-4V) into these pipelines to achieve 10 representative mutations, and the evaluated datasets cover general image classification [31], fine-grained dog breed identification [32], face recognition [33], and autonomous driving [34].
Our study reveals that MLLM-enabled mutations and traditional mutations are complementary in VDL testing: MLLMs cannot unify traditional mutations (which primarily edit the status of semantics that are already in the seed images) and exhibit low faithfulness and alignment when implementing them, e.g., none of our evaluated SoTA MLLMs can achieve image rotation. However, MLLMs bring a new dimension to image mutation: they perform impressively well on “semantic-replacement” mutations that replace semantics in the seed images (e.g., “change the background to a library”); these are uniquely enabled by MLLMs. Besides, our results show that the two mutation pipelines of MLLMs may fit distinct scenarios, and call for careful calibrations when using them. Moreover, we find that existing validation metrics are less applicable to MLLM-based mutations due to the diversified mutants and finer-grained operations, and suggest potential improvements. In sum, this paper makes the following key contributions:
  • This paper conducts the first study on the quality of test inputs generated by MLLM for VDL testing, and designs a large-scale human evaluation from the validity, alignment, and faithfulness perspectives.
  • Our results reveal the (in-)capabilities of MLLMs in image mutation, the limitation of existing validation metrics, and the distinct suitability of different mutation pipelines that are designed for MLLMs. We also summarize the lessons learned and provide suggestions for future improvements.
  • Our findings better position MLLMs and their enabled instruction-driven mutations in VDL testing. While MLLMs cannot edit existing semantics in images well like traditional mutations, they enable replacing semantics with new ones, being complementary to traditional mutations and expanding VDL testing’s scope.
2 Background: VDL Testing
In general, a visual deep learning (VDL) system typically consists of a visual understanding module and a decision module. The visual understanding module takes an image as input and extracts features from its content. The decision module yields outputs (according to the specific VDL application domain) based on the extracted features. From this perspective, image classifiers, object detectors, image captioners, and auto-driving systems, etc., that were widely tested in our community can all be counted as VDL systems [1, 2, 3, 4, 5].
The vanilla way of evaluating VDL systems is comparing their outputs with human annotations over a set of test inputs (i.e., labeled images in standard test datasets). However, manually annotating test inputs is labor-intensive, impeding automated and large-scale evaluations. Software testing has achieved remarkable success in evaluating the reliability of VDL systems. It uses different input mutation schemes to generate diverse test inputs and forms oracles to enable automatic “annotations” for the test inputs. For example, a slightly rescaled image (e.g., no less than $50\%$, as determined by the validation metric) should have the same output as the original image when fed into an object detector. This way, software testing can identify VDL faults as violations of testing oracles, where human annotations are no longer required.
Test Input Generation and Validation. Since VDL processes the semantics of image contents, various mutations and validation metrics have been proposed to generate legitimate test inputs (see details in Sec. 3). In short, the mutations broadly modify different semantics to explore potential application scenarios of VDL systems. The numeric validation metrics quantify the extents of achieved mutations and the validity of test inputs, in order to control the mutation and ensure meaningful testing results. Note that different from adversarial perturbations [^2] Such perturbations are leveraged by adversarial attacks to generate malicious inputs (i.e., adversarial examples) to fool a VDL system [35]. that add invisible noise to images and study the “worst-case” attack surfaces of VDL systems, mutations employed in VDL testing holistically alter images to diversify image semantics, exploring different usage scenarios of VDL systems.
Testing Oracle. Metamorphic testing (MT) stands out as one mainstream testing paradigm of VDL testing in recent years [2, 3, 5, 1, 36, 6, 7]. MT checks whether a VDL system’s outputs are aligned with the relation (e.g., an equality relation) defined in an MT testing oracle, when the VDL is processing an input and its mutated version. Unalignment shows that one of the inputs is ill-processed and a VDL fault is triggered. Since testing oracles are critical to detect VDL faults, input mutations themselves should properly preserve the oracles (e.g., within specific semantic ranges, as determined by the validation metric); see examples in Sec. 3.
3 Image Mutations and Validation
Figure 1: Decomposition of image semantics and their corresponding mutation schemes.
Image Decomposition: A Motivating Example. To understand VDL and the design of input mutations, we first decompose the dog image in Fig. 1 into different semantics. Starting from a lower level, the image is a matrix of pixels, and the magnitude of pixel values determines the brightness and contrast of the image. Besides, the spatial arrangement of pixels reflects geometrical properties of the image, such as the scale and the rotation angle. A reliable VDL system that recognizes dogs should be resilient to changes of these semantics. At a higher level, the image contains a dog and the background grass; the dog itself has rich semantics such as the action and the orientation. Similarly, the semantics in background may vary with the working scenario of the VDL, e.g., the grass may be replaced by snow in a winter scene. Overall, the dog-recognition VDL should focus on the semantics of the dog and be invariant to irrelevant semantics in the background.
Based on the above semantic decomposition, the following image mutation schemes have been proposed by existing works.
3.1 Explicit & Mathematical Transformations
Pixel-level. Mutations belonging to this category can be formed as uniformly changing pixel values via $\mathbf{x}$ op $c$, where $c$ is a scalar value and $\mathbf{x}$ is a matrix denoting the seed image. op decides the mutated semantics. For example, op is the add operation if the mutation changes the image brightness, i.e., adding all pixel values in $\mathbf{x}$ with a positive number $c$ increases $\mathbf{x}$’s brightness. Similarly, mutating image contrast can be formulated as $\mathbf{x}\times c$, which enlarges the difference between pixel values in $\mathbf{x}$. Representative examples are shown in Fig. 1.
Geometrical-level. As geometrical properties are determined by the spatial arrangement of pixels, existing works implement these mutations using affine transformations, which essentially move pixels to new localizations and preserve the image’s parallelism (i.e., two parallel lines remain parallel in the mutated image). Formally, suppose pixels of the image $\mathbf{x}$ are in a rectangular coordinate and the centering pixel is the origin, the $(i,j)$-th pixel’s localization in the mutated image can be calculated as $[i^{\prime},j^{\prime}]^{\intercal}=\mathbf{A}\times[i,j]^{\intercal}+\mathbf{b}$, where $\mathbf{A}$ and $\mathbf{b}$ characterize the geometrical mutation. For example, $(\mathbf{A}_{S},\mathbf{b}_{S})$ and $(\mathbf{A}_{T},\mathbf{b}_{T})$, w.r.t. scaling and translation mutations, are shown in Eq. 1,
$\mathbf{A}_{R}=\begin{bmatrix}\cos\theta,&-\sin\theta\\ \sin\theta,&\cos\theta\\ \end{bmatrix},\mathbf{b}_{R}=\mathbf{0};\;\mathbf{A}_{T}=\mathbf{0},\mathbf{b}_{T}=\begin{bmatrix}t_{x}\\ t_{y}\\ \end{bmatrix}$ (1)
where $\theta$ specifies the rotating angle. $t_{x}$ and $t_{y}$ indicate the vertical and horizontal translation distances (see Fig. 1).
Validation and Application Scope. Validating the above mutations is straightforward: developers can directly control the mutations via their parameters (e.g., $c$ in pixel-level mutations) and ensure generating valid test inputs by setting reasonable parameter values. These mutations are mostly adopted in testing safety critical VDL systems like autonomous driving [4, 5] due to their neat and explicit formulations. In addition, given these mutations’ applicability to almost all RGB images, they are widely used in testing more general VDL tasks like image classification [1].
3.2 Implicit & Data Exploration
Recent works diversify the mutated semantics with data-driven mutations. These mutations implicitly explore patterns in some images and then apply the patterns to mutate others.
Style-level & Application Scope. From a high level, each image can be decomposed as its style and content, and VDL systems should primarily focus on the content rather than style. Thus, existing VDL testing works propose style-level mutations, which generate images of distinct styles but presumably retain the content of the seed image. The “style” has been studied from two aspects. First, the style can be viewed as describing how the image is produced, e.g., the seed image in Fig. 1 is taken in reality and its style is “natural”. In that sense, the style can be changed by generating an image’s artistic-stylized variants (e.g., an image having the same dog but painted using a Van Gogh style, as in Fig. 1). These mutations over “artistic styles” have been shown effective in flipping image classifier’s outputs in recent testing works  [37, 7], and some works also employ them to reveal VDL system’s bias to non-content semantics [38], showing their high value in VDL research.
Besides, some other works decompose the weather condition as the style and separate it from the main content. The weather conditions are transferred between images to generate test inputs, as shown in Fig. 1. However, these weather-targeted style-level mutations are specifically designed for testing VDL-based autonomous driving [6], and show limited applications in other testing scenarios.
Perceptual-level & Application Scope. Recent VDL testing focuses on fine-grained semantics of the objects in images and implements perceptual-level mutations [39, 40, 7], e.g., in Fig. 1, the dog’s eyes and mouth can be mutated to evaluate whether the VDL accurately recognizes different dogs [39]. Similar techniques can be applied to face photos for testing and improving finer-grained VDL tasks like face recognition [41, 42], face aging [43, 44], crowd counting [45, 46], etc.
Validation. Both style-level and perceptual-level mutations are implemented based on generative models like GANs [47, 6, 33]. To control the mutations, existing works propose optimization-based solutions, e.g., the cycle-consistency mechanism is formed during training in [47, 6] to better separate styles from images. However, such constraints derived from optimizations lack explicit guarantee, e.g., the cycle-consistency can be violated when transferring styles among images of dissimilar objects [48, 7]. Hence, existing testing works additionally adopt numeric metrics, such as Inception Score (IS) [49], Frechet Inception Distance (FID) [50], to measure the validity of generated test inputs. Since variations on fine-grained semantics in style- and perceptual-level mutations can hardly be captured via pixel values, recent works propose to measure the extents of achieved mutations via image embedding [^3] A representation of the image that captures its semantics. The term embedding may be alternatively dubbed as “feature” in some works. [51, 40, 7] (e.g., obtaining style and content embeddings, or using embedding models like CLIP [15, 51] to compare image semantics with text descriptions).
3.3 MLLM-Based Mutations
In line with past image mutations in VDL testing, we explain how MLLMs can boost the mutation schemes from two aspects.
Unification & Generalization: Instruction-Driven. MLLMs have revolutionized image mutation by enabling text-to-image translation. Users can describe the desired mutation for a seed image using text instructions, and MLLMs generate the mutated image accordingly. This instruction-driven paradigm enables VDL developers to uniformly implement previous mutations introduced in Sec. 3.1 and Sec. 3.2 and apply domain-specific mutations in Sec. 3.2 to more general scenarios, by designing the proper text instruction.
Expansion: Semantic-Replacement. The instruction-driven paradigm also largely expands the covered semantics in image mutations. Unlike previous mutations that edit the status of semantics already presented in the seed image, MLLMs bring new mutations that can replace image semantics with new ones; we refer to this new category as semantic-replacement in the rest of this paper. For example, the dog image’s main body can be replaced with clothes (see Fig. 1) to further challenge the VDL system’s recognition.
Comparison with Object Insertion. Knowledgeable readers may wonder the differences between semantic-replacement mutations with the object insertion adopted in testing object-based VDL systems [2, 36, 3]. We clarify that semantic-replacement mutations operate in semantic-level and are finer-grained, which often require reasoning the relation and hierarchy between semantics (either intra- or inter-object). For example, the “dressing with clothes” mutation in Fig. 1 needs to understand and identify the dog’s body and replace it with appropriate clothes. Object insertion, in contrast, does not perceive image semantics and only relies on object locations in the image, so that the inserted object will not overlap with existing ones. Moreover, prior object insertion mutation is not end-to-end, as they rely on object localization results from object detectors; we thus omit it in this study.
4 Research Motivations
Limitations of Past Methods. Although the fast development of input mutations shows promising results in VDL testing, maintaining and developing them often requires expertise in different domains. Moreover, the more advanced style- and perceptual-level mutations show limited application scopes. For instance, weather-targeted mutations may perform worse in non-driving-scene images, and the face-related perceptual-level mutations only apply to highly aligned face photos (as they rely on a unified mask to pinpoint eyes, mouth, etc.). Some perceptual-level mutations may require manually annotating the to-be-mutated semantics [9, 13].
From these perspectives, we envision that MLLMs can offer an instruction-driven mutation paradigm, which is more general, automated, and highly flexible for VDL testing. MLLMs have shown high capabilities in natural language understanding and image synthesis, and therefore, various kinds of mutations that were hardly or tediously achieved by traditional methods may be easily implemented via MLLMs by directly providing instructions. Moreover, since various (subtle or complex) mutations are unifiedly described in natural language (as shown in Fig. 1), the mutations can be easily understood, maintained, calibrated, and extended by VDL developers and the community. [^4] In some sense, this new instruction-driven mutation paradigm is comparable to how query-based software security testing tools, e.g., CodeQL [52], maintain varying types of vulnerabilities in a unified query language.
Unclear MLLM Internals. Despite the optimism, MLLM’s image generation is hard to interpret due to the unclear internals. Per recent reports, MLLMs often fail to mutate images according to the given instructions [53], and may generate broken and ill-formed images from time to time [27, 28, 29]. Hence, we believe our community still lacks understanding of how exactly MLLMs can boost VDL testing, and a comprehensive study is urgent for VDL developers to understand MLLM’s (in-)capability of mutating images. In addition, considering that input validation is crucial to ensure meaningful testing results, it is also demanding to evaluate whether existing validation metrics can still be applied to MLLM-based mutations.
5 Research Study Setup
Our study answers the following research questions (RQs).
  • RQ1: Can MLLMs unify different traditional input mutations? We answer this RQ by assessing the quality of 1) how MLLMs re-implement traditional mutations and 2) how MLLMs extend prior domain-specific mutations to general scenarios.
  • RQ2: What benefits can MLLMs bring to input mutations in VDL testing? We answer this RQ by 1) assessing the quality of test inputs generated by the semantic-replacement mutations, and 2) evaluating their effectiveness in identifying VDL faults.
  • RQ3: Are existing input validation metrics reliable when facing MLLM-based mutations? We answer this RQ by evaluating whether existing validation metrics can reflect 1) the validity of test inputs generated by MLLMs and 2) the varied extents of different MLLM-based mutations.
Our RQs are based on the quality assessment of test inputs generated via different mutations, which requires perceiving and understanding image semantics. We therefore conduct a large-scale human study to evaluate image quality. Below, we introduce our evaluated aspects for the quality assessment.
5.1 Evaluated Aspects
Following design principles in existing input mutations and validation metrics, we consider the following three aspects.
Alignment. The achieved mutation must be aligned to the expectation. This is important as users need to know which mutated semantics trigger VDL faults and diagnose the VDL’s faulty behavior using the testing results. For example, if a rotation mutation does not rotate the image as expected, users will falsely interpret the tested VDL system as behaving correctly (false negatives).
Faithfulness. A mutation should faithfully preserve the semantics that are ought to be unchanged. For example, to close eyes of face photos, the mutation should let other semantics (e.g., background, hair, nose) remain unchanged and only close the eyes. Otherwise, the unanticipated mutation may break the testing oracle and bring false positives. The follow-up diagnosis for VDL’s faulty behaviors can also be misled.
Validity. Besides the alignment and faithfulness that focus on the mutated and unmutated semantics, we also holistically consider the validity of all semantics in the whole image. This is crucial as an aligned and faithful mutation may still generate unrealistic images. For example, as will be shown in Fig. 4, GPT-4V replaces a portrait’s eyes with closed ones (potentially from other portraits in its “database”) when asked to close the eyes. Although the mutation aligns the requirement and other semantics are untouched, the sharp transition around eyes makes the mutated image unrealistic, which rarely appear in real-world scenarios.
5.2 MLLMs and Their Pipelines
MLLMs. To date, two pipelines are proposed by existing works to employ MLLMs for image mutations. As illustrated in Fig. 2(b)-(c), these pipelines are optimized from different angles to avoid issues in prompt engineering or image generation modules, providing out-of-the-box usages for non-expert in MLLMs. In particular, the first pipeline (Fig. 2(b); see more details below) relies on a specifically fine-tuned MLLM, we therefore use the MLLM provided by the authors [54], along with this pipeline, as a whole for evaluations and refer to it as InstP2P. The second pipeline is post-hoc and supports incorporating different MLLMs; thus, we use three SoTA and representative MLLMs, including LLaVA [55], CogVLM [56], and GPT-4V [17] into this pipeline during evaluations.
Figure 2: Pipelines of MLLM-based mutations. Fig. 2(a) shows the most straightforward dialog-based pipeline. Fig. 2(b)-(c) illustrates two pipelines specifically optimized for image mutations. The unidirectional pipeline shown in Fig. 2(b) requires fine-tuning a MLLM and should be used along with this tailored MLLM. The bidirectional pipeline presented in Fig. 2(c) is post-hoc and supports incorporating different MLLMs.
Dialog-Based. The most straightforward way to mutate images is building a dialog with the MLLM, as shown in Fig. 2(a). For example, the users first provide the seed image and a prompt of the mutation instruction to GPT-4V and let GPT-4V return the mutated image. However, our preliminary explorations show that this dialog-based pipeline fails to generate test inputs in all cases we try: GPT-4V either generates an entirely new image (nearly all semantics are changed) or responds that it cannot generate the image. Similar issues are also noted in [57, 58, 59],
To alleviate this issue, one may expect to use prompt engineering to better guide the MLLM. Nevertheless, our preliminary study shows that it is hard to automatically tune MLLM prompts and achieve improved performance constantly on different images. Moreover, we argue that “tuning prompts” is less practical in our context, as VDL developers (main users of VDL testing tools) are often not experts in prompt engineering; MLLM should provide out-of-the-box support for implementing mutations. In this regard, recent works propose to equip MLLMs with Diffusion model [60], a more sophisticated image generation model, and design two representative optimization pipelines to orchestrate their specialties in cross-modal translation and image generation [57, 58, 61, 62, 59].
Unidirectional & Fine-tuning-based. As shown in the bottom of Fig. 2(b), InstP2P designs an unidirectional pipeline which directly uses a free-form text instruction and the seed image to generate the mutated image. In general, since the text instruction primarily describes the mutation and does not include details in the images, InstP2P requires fine-tuning a Diffusion model to support diverse mutations.
Specifically, as illustrated in the top and middle of Fig. 2(b), a GPT model is first employed to edit a given textual caption to obtain a new caption. Then, the two captions are fed into a pretrained a Stable Diffusion model [60] to generate two images, $i_{s}$ and $i_{t}$, respectively. Finally, a free-form textual instruction is generated using GPT to briefly summarize the changes between the two captions; this textual instruction, together with the two generated images $i_{s}$ and $i_{t}$, forms a data point for fine-tuning the Diffusion model, and the Diffusion model is tuned to generate $i_{s}$ when given $i_{t}$ and the textual instruction. After fine-tuning, the Diffusion model is able to generate a mutated image for any unseen images and textual instructions. We refer interested readers to [54] for the fine-tuning details.
Bidirectional & Post-hoc. Another pipeline, as designed in [57, 58, 61, 62, 59], leverages MLLMs to transform the mutation instruction based on the seed image, such that details in the seed image can be included in the transformed instruction. As shown in Fig. 2(c), this pipeline is categorized as bidirectional, since it implements the workflow of image $\rightarrow$ text $\rightarrow$ image. Specifically, due to MLLM’s promising capability of understanding images, users first feed the seed image to the MLLM and let the MLLM describe what the image would look like after applying a mutation. Here, the MLLM’s output is a textual description of semantic-level details in the mutated image. This finer-grained text description, together with the seed image, is then fed to the Diffusion model to generate the mutated image. The superiority of this pipeline (compared with the dialog-based one) has been empirically justified in [57, 58, 61, 59].
Fig. 3 shows examples of textual descriptions in the bidirectional pipeline. The text can precisely describe the mutation expected by the instruction. For instance, in the first case shown in Fig. 3, the description “The lizard would be surrounded by a snowy environment…” precisely captures the semantic transformation specified by the mutation instruction “What would it look like if it were snowing”, facilitating accurate execution of the intended mutation. In addition, the text includes many detailed and informative descriptions, e.g., “The snow would also cover the tree branch and the surrounding area…” in the first case, which gives more information of the mutated image to ease the mutant generation.
Figure 3: Textual descriptions generated in the bidirectional & post-hoc pipeline of MLLM-based mutations.
Overall, the bidirectional pipeline, to some extent, achieves automated prompt engineering by enriching the mutation instruction with the seed image’s details. The unidirectional pipeline, on the other hand, fine-tuning the Diffusion model to strengthen its capability in achieving image mutations. In our evaluations, we incorporate three MLLMs, namely, CogVLM, LLaVA, and GPT-4V, into the bidirectional pipeline, resulting in three MLLM-based mutation methods. We use the unidirectional pipeline with its tailored MLLM and denote this mutation method as InstP2P. It’s worth noting that CogVLM, LLaVA, GPT-4V, and InstP2P (before fine-tuning) use the same Stable Diffusion model [60] as the image generation module, ensuring a fair comparison for the capabilities of different MLLMs.
5.3 Datasets and Mutations
Table I lists the evaluated datasets. They are large-scale and representative to cover different application scenarios. ImageNet consists of 1,000 classes of real-life images, and is widely used for general image classification tasks. Dog-Breed dataset contains dog images with 120 different breeds and is a subset of ImageNet. It is constructed to improve VDL model’s accuracy in domain-specific applications. FFHQ includes human face photos; it is widely used for face recognition and annotation tasks. CityScape is composed of urban street scene images, which are popular in autonomous driving, object detection, and instance segmentation, etc.
TABLE I: Evaluated datasets.
Dataset #Classes Description VDL Tasks
ImageNet [31] 1,000 Real-life images Image Classification
Dog-Breed [32] 120 Dog images Image Classification
FFHQ [33] N/A Face photos Face recognition, annotation, etc.
CityScape [34] N/A Urban street scene images Autonomous driving, object detection, instance segmentation, etc.
TABLE II: Mutations and their text instructions in MLLMs.
Mutation Type Instruction
① Contrast Pixel Increase the contrast of the image.
② Rotation Geo. Rotate the image 180 degrees.
③ Artifying Style Change the style of this image to Van Gogh style.
④ SnowG What would it look like if it were snowing?”(for general images)
⑤ SnowD What would it look like if it were snowing?”(for driving scene photos)
⑥ Eyes Percep. Make the eyes of the person close.
⑦ Background MLLM- enabled Change the background to a library.
⑧ Clothes Dress the dog with clothes.
⑨ Legs Change the dog’s legs to be bionic.
⑩ Glasses Put on a pair of glasses for the dog.
  • SnowG: weather transfer mutations applied on general images. SnowD: weather transfer mutations applied on driving scene photos.
Traditional. We consider five representative mutations for the five categories of traditional mutations introduced in Sec. 3. Additionally, since the weather transfer mutation is affected by the to-be-mutated image, we evaluate it on both general images (SnowG in Table II) and driving scene photos (SnowD in Table II) from ImageNet and CityScape, respectively. The perceptual-level mutation is achieved using GANs [13]; it is dataset-specific and only applies to aligned face photos.
MLLM-Enabled. For the semantic-replacement mutations uniquely enabled by MLLMs, we consider four different variants operating in different granularities and manifesting varying characteristics. Background only edits background semantics, while the remaining three focus on finer-grained semantics in the main subject. Glasses is more “silent”: it replaces the original eyes with glasses + new eyes, and the new eyes should be mostly similar to the original ones, with only slight updates on the lighting condition for coherence. Clothes and Legs, in contrast, are more active by completely replacing target semantics with new ones, and their target semantics are of interest to different VDL models when recognizing the dog. Since these four mutations require specifying the mutated object in the images, to evaluate them uniformly on diverse images, we use the Dog-Breed dataset. That said, extending them to other objects/images should be straightforward: users only need to replace the word “dog” with the target one.
Prompt. To implement a mutation (either traditional or MLLM-enabled) using MLLMs, a text instruction (i.e., prompt) is required. We list the text instruction of each mutation in Table II. Recall as mentioned in Sec. 5.2, we consider two optimized MLLM pipelines that are specifically designed for achieving image mutations, so that a mutation’s prompt can be a single-sentence instruction, enabling an out-of-the-box usage. Due to such a simple form, we find that the MLLM is mostly robust to changes over those instructions, and we observe that, the instruction’s wording, grammatical structure, etc., does not notably impact the achieved mutation. We thus use the same instructions for both MLLM pipelines in our evaluations; this setup also allows us to evaluate the specialties of different pipelines and motivates future pipeline design.
TABLE III: Results of human evaluations. We mark results within $[0,1)$, $(2,4]$, and $(4,5]$ in red, blue, and green, respectively.
Metric Contrast Rotation Artifying SnowG SnowD Eyes Background Clothes Legs Glasses
Alignment LLaVA 0.8 0.0 3.2 2.3 4.3 1.3 1.3 1.9 1.8 3.8
CogVLM 0.6 0.1 3.1 2.0 3.3 0.7 1.1 1.4 1.3 2.4
GPT-4V 0.7 0.1 3.1 2.1 3.8 1.8 1.3 2.3 2.1 3.2
InstP2P 1.2 0.0 2.5 0.6 0.1 0.1 2.6 0.7 1.1 4.2
Traditional 4.9 4.9 1.7 0.1 4.2 4.1 N/A N/A N/A N/A
Faithfulness LLaVA 3.0 3.6 3.3 2.5 2.0 3.3 2.9 3.0 3.0 3.5
CogVLM 2.7 3.5 3.0 2.3 2.0 3.0 2.8 2.5 3.1 3.3
GPT-4V 3.0 3.7 3.3 2.7 2.1 3.0 3.0 2.8 3.1 3.5
InstP2P 4.9 1.0 4.2 4.6 5.0 3.3 3.0 4.2 4.5 4.3
Traditional 5.0 4.9 4.6 4.5 0.2 4.4 N/A N/A N/A N/A
Validity LLaVA 4.0 4.5 4.1 4.0 2.8 3.1 3.1 3.3 3.6 4.4
CogVLM 3.6 4.3 3.7 3.8 2.7 2.5 2.9 3.2 3.6 4.0
GPT-4V 4.0 4.2 3.9 3.8 2.9 2.5 3.1 3.3 3.8 4.5
InstP2P 4.9 0.9 4.1 4.8 4.9 3.6 2.6 4.2 4.5 4.2
Traditional 5.0 4.9 4.4 4.5 2.5 4.4 N/A N/A N/A N/A
6 Human Studies Setup
We form a group of 20 participants to evaluate the ten mutations in Table II according to the alignment, faithfulness, and validity requirements (discussed in Sec. 5.1). Since evaluating if a mutation fulfills these requirements needs inspecting how the changed semantics (both anticipated and unanticipated) may affect the downstream VDL testing and analysis, we hire 20 Ph.D. students experienced in software testing and VDL systems in our human study. For each mutation in Table II, we prepare 50 seed images to mutate. That is, for the six traditional mutations, each of them has $50\times 5=250$ mutated images, and for the four MLLM-enabled mutations, each of them has $50\times 4=200$ mutated images.
The common practice, as in most existing MLLM works [63, 64], is letting humans rank different mutated images. However, we observed that MLLMs can fail in implementing a mutation or generating valid images, which cannot be reflected in the ranking results. To address this, we design the questions as follows. Each question in our human study is formed by a seed image and one of its mutated variants (using different implementations), and a participant is asked to give a score (0 to 5) for the alignment, faithfulness, and validity of a mutated image. A zero score indicates that the mutated image fails to satisfy one requirement. For example, if the rotation mutation implemented using a MLLM does not rotate an image, the alignment score should be zero. We use scores 1-5 to assess each requirement because one mutation has maximal five different implementations (i.e., one traditional method + four MLLMs); the five different choices should be distinguishable for different implementations.
We let each participant evaluate one of the ten mutations, as interchangeably evaluating different mutations may confuse the participants, especially for subtle mutations that require careful comparison. Thus, each participant answers 200 (for MLLM-enabled mutations) or 250 (for traditional mutations) questions. To reduce personal preferences, each mutation is assigned to two different participants. The orders of the questions are shuffled to avoid potential bias due to question orders.
To rule out invalid evaluations (e.g., a participant misunderstands the requirement), following previous work [65], we insert sanity-check questions into each evaluation without notifying the participants. Specifically, we insert 10 questions where the “mutated” image is identical to the seed image, and the participant is expected to assign 0, 5, 5 scores for the alignment, faithfulness, and validity requirements, respectively. If a participant fails to pass 80% of the sanity-check questions, we will discard all his/her answered questions. It’s worth noting that all our participants pass the sanity check.
Before starting the human study, we prepare a 30-minute teaching for each participant. The teaching primarily aims to introduce the three requirements (i.e., alignment, faithfulness, and validity) with different successful and failed examples. It also gives participants suggestions on giving scores. For instance, zero indicates an unsatisfied requirement, and the scores should be distinguishable. Our human study is conducted using the Amazon Mechanical Turk platform [30]. We do not set a time limit for each question because we expect participants to carefully assess the mutated images. The participant can pause and resume the evaluation at any time to avoid fatigue. The average time to finish all evaluation questions (i.e., 200 or 250) is around two hours as reported by participants.
7 Results and Findings.
This section presents the results of our human study and answers three RQs mentioned in Sec. 5.
Figure 4: Mutation examples achieved via traditional method and different MLLMs.
7.1 RQ1: Unifying and Generalizing Mutations
Table III reports the average score over all 200 $\times$ 2 or 250 $\times$ 2 samples for each mutation. Since a zero score indicates that a mutated image fails to satisfy one requirement, we deem scores within $[0,1)$ as failing and mark them in red. We view scores larger than two as satisfying the requirement. Specifically, scores within $(2,4]$ are marked in blue, whereas scores within $(4,5]$ are marked in green to highlight the best method for one mutation.
Fig. 4 shows mutation examples of different methods. Below, we interpret our human study results with these examples.
Contrast and Rotation. Our results show that MLLMs are unable to implement pixel- and geometrical-level mutations, as supported by the red scores under the alignment evaluation. Note that pixel- and geometrical-level mutations directly operate on image pixels, whereas the remaining mutations focus on image features (i.e., image representations obtained using neural networks). MLLMs also primarily accept image features, leaving lower-level image semantics untouched. This explains why MLLMs fail in implementing pixel- and geometrical-level mutations, as images before and after these mutations often do not lead to notable changes in image features. Recent MLLM papers have similar observations: MLLMs perform worse in the spatial reasoning of image contents [66, 67, 68].
Different from others, InstP2P fails in the validity evaluation of Rotation and the faithfulness evaluation also has an average score of one. As shown in Fig. 4, InstP2P generates a meaningless image when implementing Rotation, and we find that similar failures occur in all evaluated samples. Recall as noted in Sec. 5.2, InstP2P differs from other MLLMs by its unidirectional pipeline: to complement the missing details in the text instruction, InstP2P fine-tunes the image generator (i.e., the Diffusion model) with mutation samples. We therefore suspect that the fine-tuning harms the generative capability of the Diffusion model. Hence, when implementing geometrical-level mutations that require spatial reasoning, InstP2P fails to generate valid images. In contrast, as other MLLMs do not modify the image generator, they mostly “transmit” the seed images if they cannot implement a mutation.
Artifying. MLLMs outperform traditional methods in transferring the artistic style of images. Past methods often extract the artistic style from one or a few target images, whose understanding of the target style may be limited. Moreover, traditional style transfer may fail if contents in the seed image and the target have distinct geometrical structures. These two limitations explain why the alignment score of the traditional method is unsatisfactory (i.e., only 1.7). MLLMs encode general knowledge and can better comprehend the target style from their immense amounts of training data.
However, MLLM’s faithfulness results are relatively worse than those of traditional methods. As shown in Fig. 4, besides transferring the style, MLLMs also edit other irrelevant semantics in the mutated images. This observation is consistent with our findings mentioned above: MLLMs often focus on high-level features and neglect low-level details. As a result, such details are lost and consequently “inpainted” with random ones when generating mutated images.
SnowG and SnowD. InstP2P fails to transfer weather conditions for general images and driving scenes. We attribute the reason to the lack of fine-tuning data. To support transferring weather conditions, the fine-tuning in InstP2P requires mutation variants of the seed image that contain different weather conditions, and such data tuples are hard to prepare on a large scale. Thus, the fine-tuned Diffusion model in InstP2P is unable to capture the weather conditions, transmitting the input image to the output in most cases, as indicated by the high faithfulness and validity score. The traditional method leverages the cycle-consistency to achieve weather transfer without requiring paired images as the guidance. This paradigm successfully transfers the weather conditions for diving scenes, resulting in a high alignment score. However, it fails to retain the faithfulness (i.e., a 0.2 score), such that mutated images do not preserve irrelevant semantics, as presented in Fig. 4. In addition, the cycle-consistency relies on the similar geometrical structure between the seed and target images; this explains why traditional method is inapplicable to mutate weather conditions in general images whose geometrical structures are largely distinct from driving scenes.
The other three MLLMs, especially GPT-4V, have satisfactory alignment, faithfulness, and validity when transferring weather conditions for both general images and driving scenes. We attribute these to the general knowledge encoded in MLLMs. Although fine-grained details are also modified in MLLM’s generated images (see Fig. 4), we envision MLLM’s promising application in testing VDL systems where weather conditions are crucial (e.g., auto-driving).
Eyes. Traditional methods outperform MLLMs when implementing perceptual-level mutations. All MLLMs are unable to deliver satisfactory alignment scores. As in Fig. 4, all MLLMs do not close the eyes but replace them with new eyes in most cases. That is, these MLLMs understand that the mutation modifying eyes but cannot precisely mutate eyes. Moreover, recall that as mentioned in Sec. 4, the traditional implementation of Eyes requires specifying the localization of eyes. To make the comparison fair, we also provide the eye locations to MLLMs. However, as shown in Fig. 4, MLLMs fail to generate smoothly mutated images and the edited eyes look unnatural, leading to relatively lower validity scores.
Answer to RQ1: MLLMs cannot implement pixel- and geometrical-level mutations well, potentially due to the lost information of such low-level semantics when MLLMs extract image features. MLLMs are also unable to generalize perceptual-level mutations. However, MLLMs optimized with the bidirectional pipeline perform better in extending style-level mutations to more general scenarios. They also exhibit better resilience: since they do not modify the image generator, they will not generate invalid images in case they cannot achieve a mutation.
7.2 RQ2: Expanding Mutated Semantics
MLLM-Enabled (⑦-⑩). As in Table III, we find that only InstP2P can implement Background to a satisfactory level, whereas the other three MLLMs tend not to edit the seed image in Background, as indicated by the relatively high faithfulness and validity score (see examples in Fig. 4). However, the trend is on the opposite when implementing Clothes and Legs, where InstP2P has lower alignment scores but much higher faithfulness and validity scores. For Glasses, all MLLMs have satisfactory alignment, faithfulness, and validity scores, and InstP2P outperforms the other three MLLMs.
Overall, these results show that MLLMs work well in semantic-replacement mutations, indicating MLLM’s high potential in expanding prior mutations with a new dimension. The varied results of different pipelines (see Sec. 5.2) reveal their specialties. Both optimized pipelines support subtle mutations to the subjects. However, the unidirectional pipeline is more suitable for background, while the bidirectional pipeline fits better when largely mutating subjects in images.
Testing Effectiveness. To further study whether MLLM-enabled mutations are useful in VDL testing, we compare their effectiveness in detecting VDL failures with traditional mutations. Contrast (①) and Rotation (②) are implemented with their original methods. Artifying (③) and SnowG (④) are implemented using MLLMs as MLLMs perform better than their original implementations. Note that SnowG denotes SnowD’s application in general images; SnowD is therefore omitted. We also omit Eyes in this evaluation since Eyes (⑥) is inapplicable to general images and MLLM also cannot generalize it according to results in Sec. 7.1.
VDL-based classification is selected as one representative VDL application. We choose three recent SoTA VDL-based classifiers, Vision Transformer (ViT) [69], SwinTransformer (SwinT) [70], and ConvNeXt [71]. These models are officially released by PyTorch and are trained with ImageNet.
[…]
Model Contrast Rotation Artifying SnowG Background Clothes Legs Glasses
ConvNeXt LLaVA N/A N/A 7.6% 4.8% 5.2% 7.7% 7.7% 2.8%
CogVLM 9.6% 4.8% 7.6% 8.6% 10.1% 4.5%
GPT4V 9.2% 3.6% 6.8% 5.7% 9.7% 1.6%
InstructP2P 6.4% 4.4% 11.2% 0.8% 1.2% 1.2%
Traditional 6.0% 0.4% 6.8% 14.4% N/A N/A N/A N/A
SwinTransformer LLaVA N/A N/A 8.0% 4.8% 5.2% 7.7% 7.3% 2.8%
CogVLM 9.6% 4.8% 7.2% 9.4% 10.9% 4.5%
GPT4V 9.2% 4.4% 7.2% 6.1% 9.3% 1.6%
InstructP2P 5.6% 4.0% 13.2% 2.0% 2.8% 1.6%
Traditional 6.8% 0.4% 8.4% 16.0% N/A N/A N/A N/A
VisionTransformer LLaVA N/A N/A 8.0% 4.8% 4.4% 6.1% 7.7% 3.2%
CogVLM 9.6% 5.2% 7.6% 9.8% 10.5% 4.9%
GPT4V 7.6% 4.4% 7.6% 5.3% 10.1% 0.8%
InstructP2P 7.2% 3.6% 12.0% 1.6% 2.0% 1.6%
Traditional 8.4% 0.4% 7.2% 16.0% N/A N/A N/A N/A
  • * N/A indicates that the mutation cannot be implemented using MLLMs.
[…]
Answer to RQ2: MLLMs are complementary to traditional mutations and expand their mutated semantics by uniquely enabling new semantic-replacement mutations. These mutations are also effective in detecting VDL faults. The unidirectional mutation pipeline of MLLMs performs better for background semantics, whereas the bidirectional pipeline is suitable for noticeable mutations on subject semantics.
7.3 RQ3: Validation Metrics
Aligned to the three aspects considered in our human study, we evaluate whether existing validation metrics can reflect the validity of test inputs generated by MLLM-enabled mutations and capture the varied extents of MLLM-enabled mutations.
7.3.1 Validity Metrics
Following existing works [72, 73], we take two popular metrics, Inception Score (IS) [49] and Frechet Inception Distance (FID) [50], to measure the validity of the mutated images. Both IS and FID use a set of real images as the reference, and compare the mutated images with these reference images to estimate how likely the mutated images are real. Lower FID values indicates better realism, whereas higher IS values indicates better realism.
TABLE V: Validity results of MLLM-enabled mutations.
LLaVA CogVLM GPT-4V InstP2P
IS 14.51 13.71 14.82 23.93
FID 88.09 89.65 88.67 93.95
As shown in Table V, IS values can roughly distinguish InstP2P from other MLLMs. While this is aligned with our human evaluation results, we expect finer-grained differences among different MLLMs can be accurately reflected. FID values, however, show contradictory results with IS values. Overall, observations in Table V indicate the inaccuracy of existing validity metrics when facing MLLM-enabled mutations. As pointed out by previous works [74, 75], FID and IS jointly consider the validity and diversity. This was not an issue previously because seed and mutated images are highly similar when applying most traditional mutations (e.g., Rotation). However, semantic-replacement mutations enabled by MLLMs bring non-determinism: multiple possible mutants can be generated when applying one mutation to a seed image. As a result, the validity metric’s results can be falsely improved by diversified mutants in MLLM-enabled mutations.
7.3.2 Embedding-Based Metrics
TABLE VI: Anticipated results of MLLM-enabled mutations.
Mutation Textual description of mutation results
Background The background of this image is a library.
Clothes A dog wearing clothes.
Legs A dog with bionic legs.
Glasses A dog with glasses.
As mentioned in Sec. 3.2, for mutations lacking explicit formulas (i.e., style-level, perceptual-level, and MLLM-enabled mutations), existing works propose to measure them via image embeddings. Following recent MLLM papers [64, 63, 76, 51], we obtain image embeddings via the SoTA embedding model CLIP [15]. The CLIP can obtain unified embeddings for image semantics and text descriptions, enabling their similarity measurement. We first prepare the textual description of the anticipated mutation result for each mutation, as shown in Table VI. For example, when mutating a dog image with Glasses, the mutated image should have “a dog with glasses”. We then use the CLIP to extract the embedding vectors for the mutated image and its corresponding textual description in Table VI; their cosine similarity is adopted to quantify the alignment of the mutation.
The faithfulness evaluation requires separating the local and fine-grained semantics such as eyes and legs from the image, which is challenging especially when the locations and postures of the dog are diverse. Therefore, we design a coarse-grained metric that only considers an image’s background and subject. Specifically, following [77], we employ the SoTA subject extractor InSPyReNet [78] to separate the subjects and background for the seed and mutated images. For Background, we use the similarity of the two subjects in the seed and mutated image as the faithfulness result, whereas for ⑧-⑩, we use the similarity of two backgrounds for faithfulness. Since semantic-replacement mutations may alter the subject’s outline (e.g., adding clothes may change the dog’s shape), we compute the similarity as the cosine similarity between CLIP image embeddings of the extracted subjects/backgrounds.
TABLE VII: Alignment (left) and faithfulness (right) results of MLLM-enabled mutations.
Background Clothes Legs Glasses
LLaVA 0.67 0.96 0.49 0.97 0.43 0.97 0.50 0.96
CogVLM 0.67 0.96 0.49 0.97 0.43 0.96 0.50 0.96
GPT-4V 0.67 0.96 0.49 0.97 0.43 0.97 0.50 0.97
InstP2P 0.69 0.95 0.50 0.97 0.44 0.97 0.52 0.97
Table VII presents the results of alignment and faithfulness evaluations for four MLLM-enabled mutations using popular embedding-based metrics. The alignment results of Background are relatively better than that of Clothes, Legs, and Glasses, indicating that the metric can detect the mutated background, but is unable to detect the mutated finer-grained semantics on subjects. The high faithfulness results show that the embedding-based metric can reflect the (ought-to-be) unchanged semantics. Nevertheless, unlike our human studies that show notable differences in alignment and faithfulness results of different MLLMs, the two numerical metrics cannot distinguish their varied results, despite that similar methods show promising results for traditional mutations [51].
Note that traditional mutations like style- and perceptual-level mutations only apply to specific domains, and accordingly, their corresponding image embeddings can be limited to the target domain to improve precision. E.g., when measuring mutations applied to face photos, existing works leverage face generators to obtain face embeddings and do not consider non-face images. Similarly, when measuring the achieved style-level mutations, prior works leverage style extractor to specifically compute the style distance. However, MLLM-enabled mutations are applicable to more general images. Although CLIP can obtain embeddings for diverse images from different domains, these embeddings consequently sacrifice their precision; similar observations are also noted in recent research [79, 80].
Answer to RQ3: Prior validation metrics are less applicable to MLLM-enabled mutations. Validity metrics like IS and FID jointly consider validity and diversity, whose numerical results can be misled by more diversified test inputs generated via MLLMs. Adapting metrics that separate validity and diversity (e.g., the Precision & Recall metric [74, 75]) for MLLMs may be a good starting point. Embedding-based metrics also become inaccurate due to the broader application scopes of MLLM-enabled mutations. Considering the high challenge of designing unified and accurate embedding models for diverse images, leveraging MLLMs to generate textual descriptions for mutants (following the bidirectional pipeline in Sec. 5.2) may be promising to measure the achieved extents of MLLM-enabled mutations.
8 Discussion and Threat To Validity
This paper aims to understand the status quo of MLLM-based VDL testing and provide insights. We also explicate the further efforts required to link MLLMs with VDL testing. Our study enhances the confidence in using MLLMs (and in general, LLMs) within the testing community. While to date, one can expect an “out-of-the-box” usage of MLLMs for testing VDL systems under some real-world scenarios, we present the following lessons and discussions.
New Mutations. Our evaluation shows that MLLMs can generalize mutations transferring semantics between images, e.g., style-level mutations that transfer styles, and enable semantic-replacement mutations that add new semantics (from other images) to the seed. These are highly valuable in VDL testing, as they can generate diverse and valid test inputs, and importantly, these mutations are largely limited or hardly feasible with traditional methods. We thus believe it is high time to incorporate MLLMs into VDL testing tasks, which shall enable comprehensive and finer-grained testing of VDL systems where subtle semantic changes are critical and can lead to catastrophic failures, such as in autonomous driving and drone navigation.
Nevertheless, contemporary MLLMs are less capable of editing existing semantics. That said, despite the optimism, MLLMs are not a panacea for all VDL testing tasks. We thus advocate to integrate and carefully choose MLLMs and traditional methods in VDL testing.
Threat to Validity. The human study may be biased, and to mitigate this threat, we have involved 20 participants with relevant expertise, and each question is evaluated by different participants to reduce personal bias. We also conduct a sanity check to rule out answers from participants who likely misunderstand the questions. We presume that it is sufficient to guarantee the credibility of our study.
Second, we also study the specialties of different pipelines of employing MLLMs for mutations. All two available pipelines are considered in our evaluations, and we incorporate four SoTA MLLMs into them to implement 10 representative mutations. We believe the evaluated cases are comprehensive to alleviate potential bias.
Third, the suitability of existing validation metrics for MLLM-based mutations are also evaluated. To make the selected metrics representative, we consider metrics that are widely adopted and have been proven effective in existing testing works. We believe this setup is sufficient to reflect the common issues in validation metrics tailored for traditional mutations.
9 Conclusion
This paper presents an in-depth study on MLLM-based mutations in VDL testing and reveals the complementary positioning of MLLMs and traditional mutations. We provide insights into the (in-)adequacy of MLLMs for VDL testing, the specialties of different mutation pipelines of MLLMs, and the limitations of existing validation metrics when facing MLLM-based mutations. Our findings call for careful integration of MLLMs and traditional image mutations for better testing effectiveness and application scenarios.